As an admirer of Thomas Young, you perform a double-slit experiment in his honor. You set your slits 1.25 mm apart and position your screen 3.83 m from the slits. While Young had to struggle to achieve a monochromatic light beam of sufficient intensity, you simply turn on a laser with a wavelength of 631 nm. How far on the screen are the first bright fringe and the second dark fringe from the central bright fringe? Express your answers in millimeters.

a) First bright fringe:

b) Second dark fringe:
Answer #1

(a) For the first bright bringe,

yB = m \lambda D / d                            { eq.1 }

where, m = finges = 1

\lambda = wavelength of the light = 631 x 10-9 m

D = distance between the slits and screen = 3.83 m

d = distance between two slits = 1.25 x 10-3 m

inserting all these values in above eq.

yB = (1) (631 x 10-9 m) (3.83 m) / (1.25 x 10-3 m)

yB = (2416.7 x 10-9 m2) / (1.25 x 10-3 m)

yB = 1933.36 x 10-6 m

yB = 1.93 mm

(b) For the second dark fringe,

yD = (m + 1/2) \lambda D / d                            { eq.2 }

where, m = fringes = 1

inserting the values in eq.2,

yD = (1 + 1/2) (631 x 10-9 m) (3.83 m) / (1.25 x 10-3 m)

yD = (3/2) (1933.36 x 10-6 m)

yD = 2900.04 x 10-6 m

yD = 2.90 mm